Wednesday 02/08/12

10:00 A – 8:45P
  • in situs day 2 (no KE)
  • antibody test try 2: 15 min fix <24 hr methanol treated embryos stained for PcG proteins.
  • much higher concentration of histone primary antibodies test stain (record)
  • Trial run of new protocol: acetic acid fix embryos: fix phase works fine.  Embryos still float in heptane methanol mix.
  • need to rinse repeatedly in methanol before moving to PBT, or residue heptane and tween react creating a opaque sticky colloid with floating embryos.  Embryos will sink in clean methanol.
  • embryos failed to devitilinize on first methanol heptane shake.
  • after cleaning methanol, resuspended in heptane and vortex until embryos sink. hopefully now embryos are properly devitilinized.
  • Reorganizing freezers: probe plasmids, archived embryos, and aliquots of salmon sperm and heparin have been moved to -20C in the basement.
  • Next up: separate probes into frequently used and rest of the library and move the library down to the basement.
  • Flip MP09 flies — flies lay really well, tons of larvae.  pupation not as successful.
  • Froze large cages of MPo8 and yw flies.
  • New cage of yw flies started.  Will need to make more plates soon.
  • Fixed issue with batchinsightm — parameter files seem to be lacking (save options from insight3).
  • Running analysis of Pc2 STORM data with low threshold (can always screen later image by photon count, so let’s get all possible ones processed to start with).
  • batchinisight still gives errors for ubx dataset.


This entry was posted in Summaries and tagged . Bookmark the permalink.